A local FM radio station broadcasts at a frequency of 96.2 MHz.

Calculate the wavelength at which it is broadcasting.
Wavelength =  meter

(1 MHz = 106 s -1)

Submit Answer
Answer #1

Answer -

Given,

Frequency = 96.2 MHz

wavelength = ?

We know that,

= C/f

where,

= wavelength

C = speed of light (3 * 108 m/s)

f = frequency

for 96.2 MHz,

= C/f

= (3 * 108 m/s) /(96.2 MHz)

= (0.03118503.118 m/s)/MHz ------------1

Also,

1 MHz = 106 s -1 -----------------2

Put 2 in 1,

= (3118503.118 m/s)/106 s -1

= 3.118503118 m or 3.12 m approx.

SO, FM IS BROADCASTING AT A WAVELENGTH OF 3.12 m approx. [ANSWER]